OpenStudy (abbycross167):

Can someone please check my work? Please? I would really appreciate it... Question~ Solve the systems using the Substitution method -3x + y = 0 3x - y = -12

OpenStudy (abbycross167):

My answer~ Solve one equation for either variable. -3x + y = 0 Add 3x to both sides. -3x+3x+y=0+3x Y=3x Second equation ~ 3x-y=-12 Rewrite the equation substituting y with 3x = 3x-3x=-12 Combine like term~ 3x-3x=0 = 0=-12 0 is not equal to -12, so that means the system of equations has no solution. Answer = no solution

OpenStudy (abbycross167):

What my teacher said was wrong~ good start....but you changed the 2nd equation The answer is correct, but I don't see how you got there.

OpenStudy (mathstudent55):

What does the teacher mean by "you changed the 2nd equation"? Did the problem really include the equation 3x - y = -12? Did you copy the second equation incorrectly? Does the teacher mean that you changed it later when you were solving?

OpenStudy (mathstudent55):

This system of two linear equations has two equations whose graphs are parallel lines. Since parallel lines never intersect, there are no points in common. That means there is no solution.

OpenStudy (mathstudent55):

Based on the equations you wrote in the post, I see nothing wrong in your solution. Your final answer is also correct.

OpenStudy (abbycross167):

I didn't copy the second equation incorrectly and I'm not sure what she meant... I copied what she said exactly

OpenStudy (abbycross167):

@mathstudent55

OpenStudy (mathstudent55):

I don't understand the teacher either. When you are solving a system of equations, and do every step correctly, like you showed, and you get to a false statement, like 0 = -12 in this case, that means there is no solution to the system of equations.

OpenStudy (abbycross167):

hmm ok, Thank you very much for helping me!!